OpenStudy (anonymous):

Find the domain, period, range, and amplitude of the cosine function. y = -6cos4x

OpenStudy (campbell_st):

well in general for cos if y = acos(bx) the amplitude is a the period is \[\frac{2\pi}{|b|}\] the range is -a to +a and the domain for cos is all real x hope this helps

OpenStudy (anonymous):

i'm still confused on the amplitude. i understand everything else. is it 6? @campbell_st

OpenStudy (campbell_st):

ok... so here is a simple explanation of the -6